OpenStudy (anonymous):

i need some major help with my algerbra 1 i dont understand it at all!!please help!! Question 2 (Multiple Choice Worth 1 points) When the quadratic equation y = (x − 3)(4x + 2) is written in standard form, what is the value of the coefficient "b?" > 1 > 4 >−10 >−6 Question 3 (Multiple Choice Worth 1 points) Choose all of the following equations that have a vertex that is the minimum point of the parabola. More than one answer may be chosen. > y = 2x2 + 2x + 8 >y = −16x2 + 2x − 4 >y = −x2 − 4x − 3 >y = x2 + 2x + 4 Question 4 (Multiple Choice Worth 1 points)

OpenStudy (anonymous):

Question 4 (Multiple Choice Worth 1 points) A quadratic equation has an x-squared term, an x term and a constant term. Sometimes Always Never -------------------------------------------------------------------------------- Question 5 (Multiple Choice Worth 1 points) Choose all of the following equations that have a graph in which the parabola opens downward. More than one answer may be chosen. y = 2x2 + 2x + 8 y = −4x2 + 8x − 19 y = 16x2 + 2x − 4 y = −x2 − 6x + 10 -------------------------------------------------------------------------------- Question 6 (Multiple Choice Worth 1 points) Find the vertex of the quadratic equation: y = 2x2 + 12x − 4. (−3, 50) (3, 50) (−3, −22) (−3, −58) -------------------------------------------------------------------------------- Question 7 (Essay Worth 4 points) Compare and contrast the two quadratic equations below. In order to receive full credit, use complete sentences to describe the following: The direction each parabola opens The vertex of each parabola y = x2 − 4x y = −2x2 + 8x − 12

OpenStudy (anonymous):

Your first question, listed as Question 2, requires a few steps: the right hand side of the equal sign, multiply the two brackets FOIL (that is the first step to get you started)